Please help to solve this Radical Expressions and Equations word problem. Here is the question: Boxing. The diagonal distance from corner to corner of a square boxing ring is 24 feet. Find the length of a side of the ring.
Answer by Roseghanezadeh(16)

You have to use the pythagorean theorem:
a^2 + b^2 = c^2
Now we have the hypotenuse "24" and since it is a square you know that all sides are equal.So "a = b"
Now we can use x instead of both of them so we can write:
x^2 + x^2 = 24^2
2(x^2) = 576
x^2 = 288
x = sqrt(288)
x = 16.97ft
You can also round it to get:
x = 17ft
So the length of the sides is 17ft.